Mass of antacid in sample: 1.5558 g

Antacid tablet: 1.6010 g

Composition of antacid:

Aluminum hydroxide: 160 mg

Magnesium carbonate: 105 mg

Protocol A: Pipet a 10.00 mL aliquot of the antacid sample solution into a 125 mL Erlenmeyer flask followed by adding about 10 mL of the bicarbonate-carbonate buffer solution (pH 10.0 ± 0.1). Transfer quantitatively a 34.96 mL aliquot of standard EDTA solutioN, 0.01000 M, to the flask using a buret. Boil gently the mixture for 5 min. on a hot plate to speed up the formation of Al-EDTA complex. Add 2 drops of Calmagite indicator and mix it well. The solution should appear pure blue in color. Back-titrate the solution with standardized zinc solution, 0.01000 M, until the color changes to purple at the endpoint. Added 34.21 ml of zinc solution.

Protocol B: Pipet a 10.00 mL aliquot of the antacid sample solution to a 125 mL of Erlenmeyer flask. Add about 10 mL of the acetate-acetic acid buffer solution (pH 5.0 ± 0.1) to mask the formation of Mg-EDTA complex. Transfer quantitatively a 25.00 mL aliquot of standard EDTA solution to the flask using a buret. Boil it gently on a hot plate for 5 min. to speed up the formation of Al-EDTA complex. Add 5 drops of xylenol orange indicator and mix well. Back-titrate the solution with a standardized zinc solution, 20.80 ml of zinc solution, until the color changes to light red at the endpoint. In this protocol slow titration will give good results.

Protocol C: Pipet a 10.00 mL aliquot of the antacid sample solution to a 125 mL of Erlenmeyer flask followed by adding about 10 mL of the bicarbonate-carbonate buffer solution (pH 10.0 ± 0.1). Add about 3 mL of triethanolamine (TEA) and swirl the mixture for 2 min. to enhance the formation of Al-triethanolamine complex and to mask the Al-EDTA complexation. Set the solution aside until the turbid solution becomes mostly clear for easy observation of endpoint. Add 5 drops of Calmagite indicator and mix well. The solution should appear wine red color in this moment.Direct-titrate the solution with standard EDTA solution until the color changes to pure blue at the endpoint. The used volume of EDTA was 5.41ml. Hence, in this protocol quick titration will give an ideal result.

What is the weight of the aluminum compound in the antacid you analyzed. Show calculations.
Answer #1

Weight of the antacid tablet (wtab): 16010 g Mass of antacid in sample: 1.5558 g Volume of the sample solution prepared (Vpow

1) 10.00 mL of the sample solution Total millimole (mmAl (10) = mmEDTA - mmZn), mmol

= 0.3496 * 0.208

=0.0727 mmol

2) 125.00 mL of the sample solution Total millimole [mmAl(125) = mmAl(10) *(125/10)], mmol

= 0.908 mmol

3) Al a in a tablet Total millimole [mmAl+Mg = mmAl (125) mmol

= 0.908 mmol